# Early-Stage Anomaly Detection: A Study of Model Performance on Complete vs. Partial Flows

### v1

@@ Abstract
This study investigates the efficacy of machine learning models, specifically Random Forest, in anomaly detection systems when trained on complete flow records and tested on partial flow data. We explore the performance disparity that arises when models are applied to incomplete data typical in real-world, real-time network environments. Our findings demonstrate a significant decline in model performance, with precision and recall dropping by up to 30% under certain conditions when models trained on complete flows are tested against partial flows. Conversely, models trained and tested on consistently complete or partial datasets maintain robustness, highlighting the importance of dataset consistency in training. The study reveals that a minimum of 7 packets in the test set is required for maintaining reliable detection rates. These results underscore the need for tailored training strategies that can effectively adapt to the dynamics of partial data, enhancing the practical applicability of anomaly detection systems in operational settings.
@@
early-stage anomaly detection, cybersecurity, complete flow information, partial flow information.
I Introduction
In the rapidly evolving field of network security, the effectiveness of machine learning (ML)-based anomaly detection systems is increasingly being tested in real-time environments. A fundamental limitation in current research is the gap between the conditions under which ML models are developed and validated, and the dynamic nature of real-world network operations. This paper addresses a critical aspect often overlooked: the disparity between models trained on complete flow records and the practical necessity of real-time anomaly detection using partial or evolving flow information.
Existing research predominantly relies on datasets generated to capture comprehensive flow records, encompassing the full lifecycle of network flows [7307098, 8543584, Kadri2024]. This approach, while detailed, does not accurately reflect the transient and incomplete nature of data in real-time network monitoring. In real-world scenarios, anomaly detection systems must often make swift decisions based on incomplete information—a stark contrast to the complete flow records used in standard dataset configurations. This discrepancy raises significant concerns about the transferability of research findings to real-time applications, where the speed and accuracy of anomaly detection are paramount.
To address this disconnect, our work evaluates how the performance of machine learning (ML) models, specifically Random Forest (RF), varies under different training and testing conditions. We experiment with models trained on complete flow records and tested on incomplete ones, as well as models trained on incomplete data and tested on the same. To simulate these scenarios, we define specific thresholds on packet counts and flow durations, assessing the effectiveness of the models with both fully and partially captured datasets.
Random Forest is selected for this investigation due to its common use in cybersecurity and anomaly detection, as demonstrated in [CICIDS2017, engelen2021, lanvin2023]. By incorporating RF into our analysis, we establish a common baseline for comparability, facilitating a more grounded evaluation of the models’ effectiveness in handling incomplete data sets in security contexts.
Our findings reveal that models trained on complete flows and tested on partial flows experience a notable decline in performance, with precision and recall dropping by up to 30% under certain conditions. Conversely, except for some isolated cases, models trained and tested consistently on either complete or partial flows maintain robustness, illustrating the critical impact of dataset consistency on model efficacy. This disparity emphasizes the challenge of transferring lab-based model accuracy to real-world, real-time detection systems, where data incompleteness can markedly impact detection capabilities. Our study also demonstrates that, for the dataset examined, at least 7 packets in the test set are required to maintain acceptable detection rates in real-time scenarios, underlining the need for model adjustments or enriched training strategies to handle partial data effectively.
The rest of this paper is organized as follows: Section II discusses the selection of thresholds for model evaluation, including an overview of the used dataset and the ML algorithm employed. Section III details the process of raw data preprocessing, flow measurement, and the production of complete and partial flow datasets. Section IV presents the performance comparison of the models when applied to complete and partial flows, focusing on precision, recall, and F1-score metrics. Section V interprets the results, explores the study’s limitations, discusses prior studies that have addressed similar challenges. The paper concludes in Section VI.
II Background
II-A Readiness Criteria for Flow Categorization
In this paper, we opt to investigate the performance of machine learning algorithms (MLA) where the flows are considered ready for categorisation based on two separate thresholds:
  • First we investigate the performance of MLAs under a packet count threshold, focusing on the quantitative aspects of network communication.
  • Subsequently, we explore the efficacy of these algorithms under a time window threshold, emphasizing the temporal dynamics of network flows.
This dual-pronged approach allows us to comprehensively assess the behavior and effectiveness of MLAs in scenarios that closely mirror real-world conditions.
By separately analyzing these two threshold mechanisms, we aim to provide a more detailed and nuanced understanding of MLAs in real-time anomaly detection. This methodology aligns with our objective to improve the applicability and accuracy of anomaly detection systems, ensuring they are well-equipped for the challenges presented by contemporary network environments.
II-B Dataset
In this research, we leverage the CICIDS-2017 dataset [CICIDS2017], an extensive collection of labeled network traffic flows. This dataset, gathered from July 3 to July 7, 2017, is detailed across multiple PCAP files, segmented by each day of the week from Monday to Friday. Specifically, Monday’s dataset exclusively features benign traffic, while the data from Tuesday to Friday encompasses a varied mix of both benign and malicious traffic patterns, including DoS/DDoS, Port Scan, Brute Force, and Infiltration events, with each day presenting distinct types of attacks.
Due to the complexity of the dataset’s structure, producing a day-wise evaluation that is both comprehensive and accessible presents considerable challenges. Therefore, our analysis is specifically concentrated on the segment of network traffic from Wednesday, notable for its wide range of Denial-of-Service (DoS) attacks. This focus allows for a detailed examination of typical DoS attacks—such as DoS Hulk, DoS GoldenEye, DoS Slowloris, and DoS Slowhttptest—alongside the Heartbleed attack.
@@ II-B Dataset
II-C ML Algorithm
In this study, we investigate the effects of incomplete flow information on network anomaly detection by employing the Random Forest algorithm, a technique widely recognized and frequently utilized in machine learning. RF, an ensemble learning method, builds numerous decision trees during the training phase and aggregates their predictions (in classification, this means taking the mode of the classes predicted by individual trees). This process not only boosts overall predictive accuracy but also helps prevent overfitting.
Opting for the RF algorithm enables us to draw on a method validated in related research [CICIDS2017, engelen2021, lanvin2023], thereby allowing us to uncover the broader implications of utilizing complete versus partial flow information. This approach facilitates the critical evaluation of the relevance and applicability of existing results within the context of real-time network security analytics.
III Methodology
III-A Raw Data Preprocessing
@@
III-B Flow Metering
For the critical task of processing the PCAP file, we utilized NFStream [AOUINI2022108719], a Python-based tool designed for efficient, flexible, and detailed data processing in network analysis. NFStream is particularly effective at converting raw network traffic traces into structured data suitable for advanced analytics. It features robust flow measurement and feature computation capabilities, underpinned by a versatile architecture. This adaptability is notably evident in its NFPlugin component [AOUINI2022108719], which allows for the integration of custom network functionalities. In our research, this adaptability was crucial, enabling the incorporation of a precise flow labeling methodology directly into the analytical framework.
Our labeling mechanism is designed to classify network flows into distinct categories based on predefined criteria that reflect the attack patterns and benign behaviors documented by the original dataset contributors [CICIDS2017]. It systematically assigns labels by evaluating each flow’s metadata, including source and destination IP addresses, port numbers, protocol types, packet payload sizes, and temporal initiation details. Additionally, this mechanism includes a feature that, upon activation, reverses the flow’s direction by swapping the source and destination parameters. This ensures accurate flow labeling, particularly in scenarios where packets from the same flow may be segmented into subflows.
III-C Preliminary Measurement
The objective of our preliminary measurement was to closely emulate the conditions of the original CICIDS-2017 dataset, aiding in the calibration of flow measurement settings for complete flow records. Accordingly, we adjusted the idle and active timeout settings in NFStream to 60 and 120 seconds, respectively, a change from the default parameters of 120 and 1800 seconds.
From the results obtained, several conclusions can be drawn:
@@ III-C Preliminary Measurement
Complementary insights supporting the observations detailed above are available as a digital artifact, accessible via [github-repo].
TABLE I: Anomaly Breakdown
Flow Type Count ZPL Count
DoS GoldenEye 7 916 870
DoS Hulk 158 027 594
DoS Slowhttptest 3 010 3 088
DoS Slowloris 5 192 1 805
Heartbleed 11
$\sum$ Attacks 174 156 6 357
BENIGN 310 381 13 580
$\sum$ Total 484 537 19 937
TABLE II: Wednesday Traffic TCP FIN/RST Breakdown
Metric Value
BENIGN FIN $>$ 2 3 008
Attack FIN $>$ 2 5 819
Total FIN $>$ 2 8 827
BENIGN RST $>$ 2 2 273
Attack RST $>$ 2 38 559
Total RST $>$ 2 40 832
III-D Producing Complete Flows
Following the insights from Section III-C, we configured NFStream to generate complete flows with these settings:
  • The initial timeout is maintained at 60 seconds, falling within the range of typical settings for Linux Kernel netfilter and IPv4 TCP-specific networking configurations.
@@ III-D Producing Complete Flows
The distribution of the refined dataset, containing complete flows generated in accordance with the aforementioned methodology, is detailed in Table III. Complementary insights supporting the decisions detailed above are available as a digital artifact, accessible via [github-repo].
TABLE III: Complete Flow Dataset Breakdown
Flow Type Count
DoS GoldenEye 7,917
DoS Hulk 158,680
DoS Slowhttptest 3,707
DoS Slowloris 5,683
$\sum$ Attacks 175,987
BENIGN 326,363
$\sum$ Total 502,350
@@
Compared to the data in Table I, Table III presents a marginally higher number of flow records for specific types. This increase in Table III, even after excluding flows with ZPL, results from our TCP flow expiration policy, which segments flows with multiple, irregular sequences of TCP FIN and RST flags into separate subflows. To mitigate any potential bias arising from such segmentation, post-processing steps are undertaken. In this phase, partial flows are matched with their complete counterparts using their six-tuple identification. This ensures that the initial segments of partial flows are accurately compared and aligned with the corresponding complete flows.
III-E Producing Partial Flows
In our partial flow measurement approach, we implemented two key mechanisms:
III-E1 Packet count
the first mechanism targets the precision of packet counts within flows: it ensures that only flows with an exact number of packets, as predefined, are retained for analysis. This strict selection criterion allows us to focus sharply on specific data exchange patterns, eliminating any flow that does not meet the exact packet count threshold. The range for this measurement was selected as $N_{pc}=\{2,3,4,\ldots,20\}$.
III-E2 Flow duration
the second mechanism deals with the duration of flows. Unlike the strict packet count approach, this mechanism allows for a degree of flexibility by retaining flows whose durations fall within a $\pm$20range of a specified target duration. This variance accommodates the dynamic nature of network traffic, ensuring that our analysis remains robust without being excessively restrictive. By applying this range, we acknowledge the natural fluctuations in flow durations while still maintaining a focus on our targeted temporal scope. The range for this measurement was selected as $N_{fd}=\{5,10,50,100,150,300,500,1000,5000,10000,15000,20000\}$ milliseconds.
In the process of producing partial flows, direct flow labeling has not been conducted. Instead, our focus was on identifying the complete flow counterparts for each partial flow using six-tuple hashes, subsequently assigning the corresponding labels from the complete flows to their partial counterparts. During this process, we also excluded any partial flows that did not meet the minimum value criteria specified in $N_{pc}$ and $N_{fd}$ for each respective dataset.
These methodologies, by specifically targeting packet count and flow duration, facilitate a nuanced analysis of the impact of partial flows on anomaly detection. This approach provides a comparative insight into how partial flow information influences the detection process relative to analyses based on complete flow information.
IV Results
IV-A Packet Count-based Evaluation

### v3

@@ Abstract
This study investigates the efficacy of machine learning models in network security threat detection through the critical lens of partial versus complete flow information, addressing a common gap between research settings and real-time operational needs. We systematically evaluate how a standard benchmark model, Random Forest, performs under varying training and testing conditions (complete/complete, partial/partial, complete/partial), quantifying the performance impact when dealing with the incomplete data typical in real-time environments. Our findings demonstrate a significant performance difference, with precision and recall dropping by up to 30% under certain conditions when models trained on complete flows are tested against partial flows. The study also reveals that, for the evaluated dataset and model, a minimum threshold around 7 packets in the test set appears necessary for maintaining reliable detection rates, providing valuable, quantified insights for developing more realistic real-time detection strategies.
@@
anomaly detection, network security, complete flow analysis, partial flow analysis, real-time detection
@@
1 Introduction
Real-time network anomaly detection is a cornerstone of modern cybersecurity, demanding systems that identify threats (malicious activity) swiftly and accurately as network traffic unfolds. Machine learning (ML) offers powerful tools for this, yet deploying ML detectors effectively faces a critical hurdle: they must frequently operate on partial or incomplete network flow information. We define this as data captured and analyzed before a flow naturally terminates, often due to the need for rapid decisions in monitoring systems. Such partial flows inherently contain less data—fewer packets over shorter durations—than the complete flows often used in research settings.
The challenge lies in bridging the gap between common research practices and operational requirements. While many influential studies in ML-based network anomaly detection develop and validate models using datasets predominantly composed of complete flows [7307098, 8543584, Kadri2024], this approach does not fully capture the dynamics of real-time systems. It is important to acknowledge that the challenge of incomplete data is not entirely unaddressed, as we discuss later further; researchers have explored aspects of processing incomplete flows, focusing on topics like flow export timing and feature availability over time [Jirsik2017, Jirsik2018, 9464011]. However, a predominant focus in many validation studies, particularly those benchmarking algorithm performance, remains on complete flows, sometimes even when underlying packet data that could simulate partial flows is available.
This persistent reliance on complete flow data for model development creates a potential disconnect. Models trained and optimized on the rich information within complete flows may exhibit significantly different, often degraded, performance (accuracy, reliability) when deployed to make predictions using the limited data in partial flows encountered in practice. This discrepancy raises valid questions about the direct applicability and transferability of findings derived solely from complete-flow analyses to real-world security operations.
This paper addresses this gap by providing a systematic investigation into the impact of using partial versus complete network flow information during both the training and testing phases of ML model development. Our primary contribution is quantifying the performance variations across distinct training/testing scenarios (complete-vs-complete, partial-vs-partial, and the critical complete-vs-partial mismatch) to understand the practical implications for deploying anomaly detection systems. While prior work has examined aspects of incomplete data, our focus is on this direct, systematic comparison across training paradigms. This highlights the practical limitations and performance trade-offs that emerge in operational environments—information crucial for security practitioners who must deploy and trust these systems.
To conduct this analysis, we employ Random Forest (RF), a supervised algorithm widely recognized and effective in cybersecurity [CICIDS2017, engelen2021, lanvin2023, Flood2024]. While unsupervised methods (e.g., Isolation Forest, Autoencoders) are common for anomaly detection, our goal is not to find the universally best detector, but rather to use a standard, well-understood supervised benchmark (RF) to specifically isolate and quantify the impact of data completeness (partial vs. complete) on model performance, a question relevant regardless of the learning paradigm. Using RF provides a robust baseline to explore how a standard ML technique contends with partial flow data under controlled conditions. We experiment with three distinct scenarios: i) models trained and tested on complete flows (baseline), ii) models trained and tested on partial flows (ideal partial detection), and iii) models trained on complete records but tested on partial ones (simulating a common deployment mismatch). Partial flows are generated using specific thresholds on packet counts and flow durations.
Our reappraisal reveals crucial insights: models trained on complete flows but tested on partial flows suffer a significant performance decline (up to 30% drop in precision/recall under certain conditions). Conversely, models trained and tested consistently (complete/complete or partial/partial) generally maintain robustness, barring isolated cases. This directly challenges the assumption that models validated on complete flows readily generalize to real-time scenarios dominated by incomplete data. It underscores the critical need for consistency between training/testing data and deployment data, highlighting how data incompleteness can markedly impact detection capabilities.
Furthermore, our study indicates that, for the dataset examined, a minimum threshold of at least 7 packets in the test set is necessary to maintain acceptable detection rates when testing on partial flows, suggesting the need for specific model adjustments or training strategies. By focusing on aligning model evaluation with the operational realities of real-time monitoring, our research aims to guide the development of more resilient and truly effective ML-based anomaly detection systems.
The rest of this paper is organized as follows: Section 2 presents the evaluation thresholds, dataset, and ML algorithm. Section 3 details our data preprocessing and flow generation approach. Section 4 analyzes model performance with complete versus partial flows. Section 5 interprets the results and discusses limitations. Section 6 examines related work. Section 7 concludes the paper.
2 Background
2.1 Readiness Criteria for Flow Categorization
In the realm of real-time anomaly detection, defining the precise moment when network flows contain sufficient information for reliable categorization poses significant challenges. In general, we distinguish between three common threshold types used to determine flow readiness for categorization:
  • Packet count: establishing a minimum packet count observed in a flow.
  • Flow size: categorizing flows based on a specified amount of transferred data (in bytes).
  • Time window: defining a fixed time interval from the start of the flow for data collection.
While these threshold-based approaches each offer unique advantages, they also bring inherent challenges in real-time anomaly detection. Different applications exhibit diverse traffic patterns; for instance, video streaming or file transfers generate large bursts of packets quickly, whereas DNS queries might involve fewer packets exchanged over a longer duration. Such variability means that flows reaching the same packet count or byte count threshold may have vastly different durations or represent different stages of communication maturity.
The time window approach, conversely, ensures a uniform time frame for decision-making but may capture insufficient data for short-lived flows or excessive, potentially redundant data for very active flows within the same window. Finding a single threshold type or value that is optimal across all types of network traffic and applications remains difficult.
A potential solution involves hybrid approaches, combining multiple criteria. However, determining the optimal balance between packet count, flow size, and duration to ensure comparable information richness across heterogeneous flows remains a complex task and an area of ongoing research.
2.2 Selecting Thresholds for Model Evaluation
In this paper, we focus our investigation on the performance of machine learning algorithms (MLAs) where flows are considered ready for categorization based on two well-defined and commonly considered threshold types:
  • First, we investigate performance under varying packet count thresholds, focusing on the quantitative aspect of communication exchange (number of interactions).
  • Subsequently, we explore efficacy under varying time window thresholds (flow duration), emphasizing the temporal dynamics.
This approach allows us to systematically assess the behavior and effectiveness of the chosen MLA in scenarios simulating different real-time decision points based on either accumulated packet interactions or elapsed time.
By separately analyzing these two threshold mechanisms, we aim to provide a more detailed understanding of how data incompleteness, defined by these common operational constraints, affects MLA performance in real-time anomaly detection. This methodology aligns with our objective to quantify the gap between idealized complete-flow analysis and practical deployment scenarios.
2.3 Dataset
In this research, we utilize the CICIDS-2017 dataset [CICIDS2017], a widely referenced collection of labeled network traffic flows designed for intrusion detection research. Generated over five days (July 3-7, 2017), the dataset includes benign traffic and a variety of simulated attacks, including DoS/DDoS, Port Scan, Brute Force, Web Attacks, Infiltration, Botnet, and Heartbleed, captured as PCAP files and pre-processed flow records. Monday’s traffic is exclusively benign, while Tuesday through Friday contain mixtures of benign and malicious activities, with attack types varying by day.
This dataset has been extensively used for evaluating anomaly detection systems [CICIDS2017, Hindy2021, lanvin2023, pekar2024evaluating], making it a suitable benchmark for comparing results and understanding how methodological choices (like using partial flows) impact outcomes. However, many prior works have relied on the provided complete flow records without explicitly investigating the performance implications when models face partial flows, which is the focus of our study. Furthermore, while CICIDS-2017 provides a valuable baseline, we acknowledge its age (data from 2017). The network traffic patterns and specific attack vectors may not fully represent the complexity of the current threat landscape, highlighting the need for future validation on more contemporary datasets.
Due to the dataset’s structure (different attacks/days) and the complexity of ensuring comparable analysis across days with differing traffic mixes, our analysis concentrates on the network traffic from Wednesday. This day includes a mix of DoS attacks (Hulk, GoldenEye, Slowloris, Slowhttptest) aimed at exhausting server resources (availability impact) and the Heartbleed attack, which targets a specific OpenSSL vulnerability to extract sensitive memory content (confidentiality impact). The anomalies detected by our model thus correspond to these specific labeled attack types present in the Wednesday data.
Crucially, the integrity of the original CICIDS-2017 flow records has been questioned due to identified labeling errors and inconsistencies [engelen2021, lanvin2023, Flood2024, pekar2024evaluating]. To mitigate risks to validity, our methodology does not rely on the pre-generated flows. Instead, we perform a careful regeneration and labeling of flow records directly from the original Wednesday PCAP file, following the attack descriptions provided by the dataset creators and incorporating corrections based on findings in [pekar2024evaluating], ensuring a more accurate ground truth for our experiments.
2.4 ML Algorithm
In this study, we investigate the effects of incomplete flow information on network anomaly detection by employing the Random Forest algorithm. RF is an ensemble learning method based on decision trees, widely recognized for its robustness and frequent use in cybersecurity applications, including intrusion detection [CICIDS2017, engelen2021, lanvin2023, pekar2024evaluating, Flood2024]. It operates by constructing multiple decision trees during training and outputting the mode of the classes (classification) or mean prediction (regression) of the individual trees. This ensemble approach generally yields higher accuracy and better resistance to overfitting compared to individual decision trees.
Our choice of RF is motivated by several factors. First, its prevalence in the literature provides a well-established benchmark, allowing for comparison and context within existing research. Second, as a supervised learning algorithm, it allows us to directly assess classification performance (detecting known labeled anomalies) based on varying levels of input data completeness. While unsupervised methods are often used for detecting novel anomalies, RF is suitable for our specific goal: to quantify how performance on known anomaly types changes when a standard, widely-used classifier is trained or tested with partial versus complete flow information.
This focused methodology, centered on a single representative MLA, allows for a clear analysis of data completeness impact. While evaluating multiple algorithms is valuable future work, this approach provides a necessary baseline understanding of the challenges posed by partial flows in operational network security contexts, potentially informing the interpretation of results from other studies and guiding future work using different models.
3 Methodology
This section details the systematic process used to generate complete and partial flow datasets from the raw packet captures and prepare them for ML analysis. Complementary insights, scripts, and detailed data supporting the observations and decisions described below are available as a digital artifact, accessible via [github-repo].
3.1 Raw Data Preprocessing
@@
3.2 Flow Metering and Labeling
For processing the preprocessed PCAP file and generating flow records, we utilized NFStream [AOUINI2022108719], a flexible Python-based network data analysis framework. NFStream converts raw traffic into structured flow data suitable for analysis and allows customization through its NFPlugin component. This adaptability was crucial for incorporating our specific flow labeling and generation logic directly into the metering process.
We implemented a custom labeling mechanism within NFStream. This mechanism assigns a label (e.g., ‘BENIGN’, ‘DoS Hulk’, ‘DoS GoldenEye’) to each generated flow based on criteria derived from the original dataset documentation [CICIDS2017], primarily using source/destination IP addresses, port numbers, and protocol associated with known attack infrastructure and time windows. This process ensures that flows are labeled according to the ground truth established for the dataset, directly addressing the malicious anomalies targeted in this study. The mechanism also handles bidirectional flows correctly. This integrated labeling approach, applied during flow generation, provides accurately labeled data tied to the specific flows generated under different timeout and export conditions.
3.3 Preliminary Measurement
To understand the characteristics of the traffic and inform the settings for generating our definitive “complete flow” dataset, we conducted a preliminary measurement run. This run used NFStream with idle and active timeout settings (60 and 120 seconds, respectively), reflecting potential parameters similar to the original CICIDS-2017 processing. This allowed us to observe phenomena, particularly those arising from flow segmentation caused by these timeouts, that could affect analysis if not properly handled.
Table I shows the resulting breakdown of labeled flow records. It separately counts records with packet payload $>0$ and those with the payload $=$ 0. The sum of these two counts for a given attack type represents the total number of flow records generated for that attack under these preliminary timeout settings. Table II details the occurrence of TCP FIN/RST flags within these records.
TABLE I: Initial Labeled Flow Record Counts
Flow Type Flows (payload $>0$) Flows (payload $=0$)
DoS GoldenEye $7916$ $870$
DoS Hulk 158 027158027158\,027158 027 $594$
DoS Slowhttptest $3010$ $3088$
DoS Slowloris $5192$ $1805$
Heartbleed $11$ 00
$\sum$ Attacks 174 156174156174\,156174 156 $6357$
BENIGN 310 381310381310\,381310 381 13 5801358013\,58013 580
$\sum$ Total 484 537484537484\,537484 537 19 9371993719\,93719 937
TABLE II: Flow Counts with $>2$ TCP FIN or RST Flags
Metric Flow Record Count
BENIGN Flows with FIN $>$ 2 $3008$
Attack Flows with FIN $>$ 2 $5819$
Total Flows with FIN $>$ 2 $8827$
BENIGN Flows with RST $>$ 2 $2273$
Attack Flows with RST $>$ 2 38 5593855938\,55938 559
Total Flows with RST $>$ 2 40 8324083240\,83240 832
Key observations from this preliminary run included:
@@ 3.3 Preliminary Measurement
These observations highlighted the need for specific flow generation parameters and filtering steps to create a cleaner dataset focused on the core characteristics of the traffic flows, rather than measurement artifacts or secondary effects.
3.4 Producing Complete Flows
Based on the insights from the preliminary measurements (Section 3.3), we configured NFStream to generate a refined dataset of “complete” flows (CF). Our goal was to capture the entirety of each relevant communication exchange while mitigating the artifacts observed previously. The following settings and filters were applied:
  • The idle timeout is maintained at 60 seconds, falling within the range of typical settings for Linux Kernel netfilter and IPv4 TCP-specific networking configurations.
@@ 3.4 Producing Complete Flows
TABLE III: Refined Complete Flow Dataset Breakdown
Flow Type Flow Count
DoS GoldenEye $7917$
DoS Hulk 158 680158680158\,680158 680
DoS Slowhttptest $3707$
DoS Slowloris $5683$
$\sum$ Attacks 175 987175987175\,987175 987
BENIGN 326 363326363326\,363326 363
$\sum$ Total 502 350502350502\,350502 350
@@
The resulting distribution of the refined CF dataset is detailed in Table III. Compared to Table I, it presents a marginally higher number of flow records for specific types. This is primarily due to our TCP flow expiration policy, which segments flows with multiple, irregular sequences of TCP FIN and RST flags into separate subflows. To mitigate any potential bias arising from such segmentation, post-processing steps are undertaken, where partial flows are matched with their complete counterparts using their six-tuple identification. This ensures that the initial segments of partial flows are accurately compared and aligned with the corresponding complete flows. This CF dataset represents our baseline for “ideal” or complete information available post-hoc.
3.5 Producing Partial Flows
To simulate real-time detection scenarios where decisions must be made before a flow completes, we generated partial flow (PF) datasets based on the CF dataset. For each complete flow in the CF dataset, we generated corresponding partial flows exported at specific thresholds using two distinct mechanisms, reflecting common real-time triggers:
  1. 1.
    Packet Count Threshold: This mechanism exports a partial flow record precisely when the flow reaches a specific number of packets (from the start of the flow). We generated separate PF datasets for exact packet counts $N_{pc}$ across the range $\{2,3,4,\ldots,20\}$. This simulates scenarios where analysis is triggered after a fixed number of packet exchanges.
  2. 2.
[…]
This methodology allows us to directly compare the performance of ML models trained/tested on the CF dataset versus various PF datasets, systematically examining the impact of information availability defined by these common real-time export triggers (packet count and duration). This contrasts with studies relying solely on pre-defined complete flows and enables a quantitative assessment relevant to operational deployment.
4 Results
This section presents the empirical evaluation designed to quantify the performance impact of using partial versus complete flow information for anomaly detection with Random Forest on the prepared CICIDS-2017 Wednesday dataset. We analyze results based on two common partial flow generation triggers: packet count and flow duration. The findings provide crucial context for interpreting previous studies that relied primarily on complete flows and highlight challenges for real-time deployment.
Evaluation Metrics: For binary classification scenarios, we report precision, recall, and F1-score for the anomaly class to focus on attack detection capability, which is the primary concern in cybersecurity applications where missing attacks has severe consequences. For multi-class scenarios, we report macro-averaged metrics across all classes. Due to these different evaluation focuses and the inherent class imbalance in network traffic, direct performance comparisons between binary and multi-class results should be interpreted with caution. For consistent cross-paradigm comparisons, we provide alternative metric calculations in our digital artifacts [github-repo].
4.1 Packet Count-based Evaluation
We first examine scenarios where partial flows are defined by reaching a specific packet count ($N_{pc}$).